OpenStudy (anonymous):

Simplify 6 – 4x – 2 + x. (Points : 4) 4 + x 3 – 4x 1 4 – 3x

OpenStudy (anonymous):

all you need is to sum and subtract the numbers

OpenStudy (anonymous):

so 6 - 4 - 2?

OpenStudy (anonymous):

Lets change it up! 4x + x = 5x = Number one 6 - 2 = 4 = Number 2 4 - 5x is the answer!

OpenStudy (anonymous):

so you have -4x + x <- combine them and then 6 - 2 <- combine them and you get your answer.

OpenStudy (anonymous):

Ok Thank you both

OpenStudy (anonymous):

No problem pss Best Response Me!

OpenStudy (anonymous):

wait that dosent make sence, because 4 - 5x isnt an answer

OpenStudy (anonymous):

separate both but they are in the same place. Imagine the numbers in front of x to be one group and those that doesn't have a x to be another group

OpenStudy (anonymous):

for example 4x-2x+5-3 that means you have 4x's minus 2x's so you get 2x and then you have +5 -3 so that is equals to 2 therefore you have 2x+2

OpenStudy (anonymous):

so it would be 4 + x

OpenStudy (anonymous):

no, how much is -4x + x? Remember that x alone means 1x

OpenStudy (anonymous):

-4x + 1 so then -4 + 1 so it would be -3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now you have -3x. Now is time for those that doesn't have x's which is +6 -2

OpenStudy (anonymous):

6 - 2 = 4

OpenStudy (anonymous):

so now you have a positive 4. Therefore you have -3x +4. Which one looks like this answer ?

OpenStudy (anonymous):

3 – 4x

OpenStudy (anonymous):

wow ok

OpenStudy (anonymous):

3 - 4x

OpenStudy (anonymous):

from where do you get a -4x? if your answer is -3x +4

OpenStudy (anonymous):

ohhhhh... it didnt occur what i did wrong, sorry, the answer is 4 - 3x... right? or am i just purely lost?

OpenStudy (anonymous):

You are correct

OpenStudy (anonymous):

Ok, Thank you

OpenStudy (anonymous):

No problem